You have just purchased a municipal bond with a $10,000 par value for $9,500. You purchased it immediately after the previous owner received a semiannual interest payment. The bond rate is 6.6% per year payable semiannually. You plan to hold the bond for 4 years, selling the bond immediately after you receive the interest payment. If your desired nominal yield is 4% per year compounded semiannually, what will be your minimum selling price for the bond?
Answer #1

ANSWER:

N = 4 YEARS OR 8 SEMI ANNUAL YEARS

BOND RATE = 6.6% OR 6.6% / 2 = 3.3% SEMI ANNUALLY

SEMI ANNUAL PAYMENT = BOND RATE * PAR VALUE = 3.3% * 10,000 = 330

PURCHASING VALUE OF BOND = $9,500

I = 4% PER YEAR OR 4% / 2 = 2% PER SEMI ANNUALLY

FV = PURCHASING VALUE OF BOND(F/P,I,N) - SEMI ANNUAL PAYMENT(F/A,I,N)

FV = 9,500(F/P,2%,8) - 330(F/A,2%,8)

FV = 9,500 * 1.172 - 330 * 8.583

FV = 11,134 - 2,832.39

FV = 8,301.61

SO THE MINIMUM SELLING PRICE OF BOND IS $8,301.61